Random sample of size 25 from a normally distributed population with the mean of 4 and variance of 16. Find the point estimate of the population mean if P(X>x)=0.9798?
1
Expert's answer
2021-09-13T17:13:56-0400
"P(X<x)=1-P(Z\leq\dfrac{x-\mu}{\sigma/\sqrt{n}})"

"=1-P(Z\leq\dfrac{x-4}{\sqrt{16}/\sqrt{25}})=0.9798"

"P(Z\leq1.25(x-4))=0.0202"

"1.25(x-4)\approx-2.04963"

"x=2.36"

"x=2.36"

The point estimate of the population mean is 4.